SCHEDULED SPEAKERS / CONFÉRENCIERS PRÉVUS

Here is a list of the scheduled speakers, along with the titles of their talks where available. Abstracts for all talks may be found at the following world wide web page:

http://www.cms.math.ca/Events/winter98/

Voici les conférenciers prévus, ainsi que les titres de leurs conférences. Les résumés pour tous les conférences seront disponibles à l’adresse Web suivante:

http://www.cms.math.ca/Events/winter98/

COXETER-JAMES LECTURE CONFÉRENCE COXETER-JAMES

Henri Darmon (McGill University) Recent progress in the theory of elliptic curves

DOCTORAL PRIZE LECTURE CONFÉRENCE DU PRIX DOCTORAL

Yuri Berest (University of California, Berkeley) Lacunae for hyperbolic differential operators with variable coefficients

PLENARY SPEAKERS CONFÉRENCIERS PRINCIPAUX

Miklos Csörgo (Carleton University) Random walking around financial mathematics

Z. Füredi (University of Illinois at Urbana / Inst. Hungarian Acad. Sci.) Lotto, football pool and other covering radius problems

W. Goldman (University of Maryland) Topology and dynamics of moduli spaces of geometric structures on surfaces

Donal O’Shea (Mt. Holyoke College) The glory and perversity of us

PUBLIC LECTURE CONFÉRENCE PUBLIQUE

H.S.M. Coxeter (University of Toronto) The Descartes circle theorem and Fibonacci numbers

SPECIAL SESSIONS / SÉANCES SPÉCIALES Algebraic Geometry / Géométrie algébrique

(Org: P. Milman)

Fedya Bogomolov (NYU Courant Institute) On the fundamental groups and universal coverings of complex projective surfaces and symplectic fourfolds

Andrew Hwang (University of Toronto) Construction of Circle-Invariant Kahler Metrics

Lisa Jeffrey (University of Toronto) The Verlinde formula for moduli spaces of parabolic bundles

Askold Khovansky (University of Toronto) Algebraic Geometry and Geometry of convex polyhedra

Donal O’Shea (Mt. Holyoke College) Limits of tangent spaces to real surfaces

Evgenii Shustin (Tel-Aviv University) Asymtotically proper bounds in the geometry of equisingular families of curves

Yosef Yomdin (Weizmann Institute) A tractable problem of the way (hopefully) to counting limit cycles of polynomial vector fields

Discrete Geometry / Géométrie discrète (Org: Robert Erdahl, Marjorie Senechal, Walter

Whiteley)

V. Alexandrov (University of Moscow) Sufficient conditions for the extendibility of an N-th order flex of polyhedra

Valery Alexeev (University of Georgia) Families of algebraic varieties associated with cell decompositions

Pau Atela (Smith College) Periodicity in geometric dynamical models in Phyllotaxis Lynn Batten (University of Manitoba) Blocking sets and security

Margaret Bayer (University of Kansas) Eulerian partially ordered sets

András Bezdek (Auburn University) A Sylvester type theorem on circles

Karoly Bezdek (University of Budapest) On a stronger form of Rogers’ lemma and the minimum surface area of Varoni cells in unit ball packings

T. Bisztriczky (University of Calgary) A signature theorem for unifrom oriented matroids

Jin-Yi Cai (University of Buffalo) A new transferenece theorem in geometry of numbers with applications to Ajtai’s connection factor

R. Connelly (Cornell University) Holes in a membrane: tension percolation

H.S.M. Coxeter (University of Toronto) Whence does an ellipse look like a circle?

Balazs Csikos (Eötvös University, Budapest) Some results around the Hadwiger-Knesar-Poulsen conjecture

Ludwig Danzer and Gerrit van Ophuysen (University of Dortmund) A species of planar triangular tilings with inflation factor \sqrt{−τ}

Robert Dawson (Saint Mary’s University) What Shape is a Loaded Die?

Michel Deza and Mikhail Shtogrin (Ecole Normale Sup., Paris) Embedding of regular tilings and star-honeycombs

Robert Erdahl (Queen’s University) Voronoi’s hypothesis on perfect domains

Ferenc Fodor (Auburn University) Large polygons in convex sets and polygons with large perimeter

Deborah Franzblau (CUNY/College of Staten Island) Generic rigidity of molecular graphs

Christopher Gold (Laval University) Voronoi methods in geomatics – the importance of the spatial model

Chaim Goodman-Strauss (University of Arkansas) Addressing and substitution tilings

Timothy Havel (Harvard University) The role of tensegrity in distance geometry

Donald Jacobs (Michigan State University) Graph rigidity: Applications of material science and proteins Gábor Kertétz (Eötvös Loránd University) Dido problem on planes of constant curvature

Wlodzimierz Kuperberg (Auburn University) Covering the cube with equal balls

Barry Monson (University of New Brunswick) Realizations of regular abstract polytypes

Konstantin Rybnikov (Queen’s University) On traces of d-stresses in the skeletons of lower dimensions of homology manifolds

Idzhad Sabitov (Moscow State University) Solution of polyhedra

Peter Schmitt (Universität Wien) The versatility of (small) sets of prototiles

I. Talata (University of Budapest) On translative coverings of a convex body with its homothetic copies of given total volume

Anke Walz (Cornell University) The Bellows Conjecture in dimension four

Walter Whiteley (York University) Constraining a spherical polyhedron with dihedral angles

EDUCATION I/ ÉDUCATION I IDENTIFYING AND OVERCOMING BARRIERS TO TEACHING AND LEARNING MATHEMATICS AT UNIVERSITY/ IDENTIFIER ET VAINCRE LES OBSTACLES À L’ENSEIGNEMENT ET L’APPRENTISSAGE DES MATHÉMATIQUES À L’UNIVERSITÉ (Org: Morris Orzech and Grace Orzech)

Ed Barbeau (University of Toronto) The teacher as coach

William Byers (Concordia University) Revealing the inner mathematician

Morris Orzech (Queen’s University) Addressing student difficulties specific to linear algebra

Tom Rishel (Cornell University) Teaching and job initiatives for graduate programs in mathematics

Martha Siegel (Towson University)

Keith Taylor (University of Saskatchewan) The math readiness project at the University of Saskatchewan

EDUCATION II/ ÉDUCATION II TEACHING MATHEMATICS – WHY WE

DO WHAT WE DO IN THE

CLASSROOM?/ L’ENSEIGNEMENT DES

MATHÉMATIQUES – POURQUOI FAIT-ON CE QU’ON FAIT EN CLASSE? (Org: William Higginson, Morris Orzech and Grace

Orzech)

George Gadanidis (Durham Board of Education) If Piaget was a math teacher — Theoretical images, classroom practice and the power of students’ minds

William Higginson (Queen’s University) Having, knowing, and being: some fundamental questions about mathematics teaching

Lynn McAlpine and Cynthia Weston (McGill University) How six outstanding math professors use reflection to improve their teaching

Pat Rogers (York University) The importance of why: encouraging students to reflect

Nathalie Sinclair (Queen’s University) Romancing powerful mathematical ideas

Extremal Combinatorics / Combinatoire

extrémale (Org: D. de Caen)

Richard Anstee (University of British Columbia) Small forbidden configurations

Jason Brown (Dalhousie University) The inducibility of complete bipartite graphs

Ralph Faudree (University of Memphis) Anti-Ramsey Colorings in several rounds

David Fisher (University of Colorado) The minimum number of triangles in a graph

Jerrold Griggs (University of South Carolina) Extremal graphs with bounded densities of small subgraphs

Penny Haxell (University of Waterloo) Integer and fractional packings in dense graphs

Felix Lazebnik (University of Delaware) On a class of algebraically defined graphs

Laszlo Székely (University of South Carolina) Some new Erdös-Ko-Rado type problems Bing Zhou (Trent University) Extremal graphs related to star chromatic number and fractional chromatic number

GRADUATE STUDENT SESSION SÉANCE POUR ÉTUDIANTS DIPLÔMÉS

(Org: David Gregory)

Leo Butler (Queen’s University) A new class of homogenous manifolds with Liouville-integrable geodesic flows

Leo Creedon (University of Alberta) Constructing free groups in modular group algebras

Mark DeFazio (York University) The behaviour of the complex zeros of the Laguerre polynomials

Jody Esmonde (McGill University) Parametric solutions to the generalized Fermat equation

Shaun Fallet (William and Mary University) Multiplicative principal minor inequalities for totally nonnegative matrices

Andrei Gagarin (University of Manitoba) Characterizations of (α, β)-polar graphs by forbidden induced subgraphs

Malcolm Harper (McGill University) A family of Euclidean rings containing Z [\sqrt{14}]

Yu-Ru Lui (Harvard University) The Turán sieve and probabilistic Galois theory

David McKinnon (University of California, Berkeley) An arithmetic Bézout theorem

Satya Mohit (Queen’s University) The ABS-Conjecture and bounds for the order of the Tate-Shafarevich group

Daniel Piché (University of Waterloo) Wavelet compression on fractal tilings

Philip Saidak (Queen’s University) On zero-free regions of the zeta function

Gregory Smith (University of California, Berkeley) Computing global extension modules

Sarah Sumner (Queen’s University) Investigating transcendence in the field of p-adic numbers

Drew Vandeth (McQuarrie university, Australia) Transcendence of the values of generalized Mahler function MATHEMATICS ON THE INTERNET MATHÉMATIQUES SUR INTERNET

(Org: June Lester, Nathalie Sinclair and Malgorzata Dubiel)

Bill Casselman (University of British Columbia) Colour, animation, interaction – the next generation of electronic articles

Stan Devitt (Waterloo Maple) MathML support for defining notation – an essential step towards live math on the web

Stan Devitt (Waterloo Maple) The Relationship between OpenMath and MathML

Nick Jackiw (KCP Technologies) JavaSketchpad: Taking Dynamic Geometry on-line

Loki Jörgenson (CECM) The CMS journals online: A study in digital publishing

Robert Miner (WebEQ) Putting math on the Web

Number Theory/ Théorie des nombres (Org: Ram Murty, Noriko Yui)

Amir Akbary (Concordia University) On the distribution values of symmetric square L-functions in the half plane Re(s) > \frac{3}{2}

Henri Darmon (McGill University) Modularity of hypergeometric abelian varieties

Chantal David (Concordia University)

Jacek Fabrykowski (University of Manitoba)

G. Frei (Laval University) On the discovery of the reciprocity law by Artin

Eyal Goren (Corcordia & McGill University, CI-CMA) Stratifications of moduli spaces and modular forms

C. Greither (University of Laval) On Brumer’s Conjecture

James Huard (Canisius College, Carleton University) An arithmetic reciprocity relation of Liouville type and applications

Hershy Kisilevsky (Concordia University) Rank of E(K) for cyclic cubic extensions K/Q

Manfred Kolster (McMaster University) Higher relative class number formulas Arne Ledet (Queen’s University) Some small 2groups as Galois groups

Claude Levesque (University of Laval) Explicit solutions of a family of Thue diophantine equations

Kumar Murty (University of Toronto) Zeros of Dedekind zeta functions in towers of fields

Werner Nowak (University of Austria) Large convex domains sometimes contain more lattice points than we would expect

Yiannis Petridis (McGill University) Zeros of the Riemann Zeta function and central values of the Lseries of holomorphic cusp forms

Gael Rémond (University of Ottawa) Theta heights and Kodaira construction

Damien Roy (University of Ottawa) Criteria of algebraic independence and approximation by hypersurfaces

Gary Walsh (University of Ottawa) Solving families of quartic equations

Hugh Williams (University of Manitoba) Computer verification of the Ankeny-Artin-Chowla conjecture for all p < 5.10^{10}

Kenneth Williams (Carleton University) Values of the Dedekind eta function at quadratic irrationalities

Operator Algebras/ Algèbre d’opérateurs (Org: James Mingo)

Berndt Brenken (University of Calgary) Endomorphisms of finite direct sums of I_{∞} factors

Ken Davidson (University of Waterloo) Principal bimodules of nest algebras

George Elliot (University of Toronto) An abstract Brown-Douglas-Fillmore absorption theorem, II

Thierry Giordano (University of Ottawa) Ergodic theory and dimension G-spaces

Daniel Kucerovsky (Fields Institute, University of Toronto) An abstract Brown-Douglas-Fillmore absorbtion theorem, I

Michael Lamoureux (Universtiy of Calgary) Crossed product algebra constructions

A. Nica (University of Waterloo) Some minimization problems for the free analogue of the Fisher information John Phillips (University of Victoria) Spectral flow and index in bounded and unbounded θ-summable Fredholm modules

Jack Spielberg (Arizona State University) A new look at C*-algebras of infinite graphs

Sam Walters (University of Northern British Columbia) K-theory on non commutative spheres arising from the Fourier automorphism

Probability Theory/ Théorie des probabilités (Org: Miklos Csörgo)

Siva Athreya (Fields Institute) Existance of positive solutions satisfying the boundary Harnack principle for a semi-linear Dirichlet problem

M. Claude Belisle (Université Laval) The hit-and-run sampler

David Brillinger (University of California, Berkeley) Some aspects of the motion of particles described by stochastic differential equations

Murray D. Burke (University of Calgary) Model checking and estimation: A large sample approach

Colleen D. Cutler (University of Waterloo) Scaling structures, chaos, and determinism in time series

Andre R. Dabrowski (University of Ottawa) A unified approach to fast teller queues and ATM

Eric Derbez (McMaster University) Generating functions and integrated super Brownian excursion (ISE)

Shui Feng (McMaster University) The behaviour of some degenerate diffusions near boundary under large deviations

Rene Ferland (Université du Québec à Montréal) Propagation of chaos: from Physics to Finance

Antonia Foldes (The College of Staten Island, CUNY) About the local time of random walk and Brownian motion

Genevieve Gauthier (École des Hautes Études Commerciales) Service de l’enseignement des methodes quantatives de gestion

Edit Gombay (University of Alberta) Correcting some limit theorems about the likelihood ratio

Lajos Horvath (University of Utah) Best approximations for bootstrapped processes with applications Gail Ivanoff (University of Ottawa) Set-Indexed Martingales

Mike Kouritzin (University of Alberta) Parabolic equations with random coefficients

Reg Kulperger (University of Western Ontario) Empirical processes and tests of independence

Brenda MacGibbon (Université du Québec à Montréal) On statistical minimax estimation and principal Eigenfuctions of the Laplacian

Neal Madras (York University) In search of faster simulations

Don L. McLeish (University of Waterloo) Estimating parameters of financial time series using highs and lows

Majid Mojirsheibani (Carleton University) Combined estimation and probabilistic classification

Bruno Remillard (Université du Québec à Trois-Rivières) Empirical Processes Based on Pseudo-Observations

Pál Révész (Math. Inst. Hung. Acad Sci.) Critical branching Weiner process

Jeffrey S. Rosenthal (University of Toronto) The mathematics of Merkov chain Monte Carlo algorithims

Tom Salisbury (York University) The complement of the planar Brownian path

Byron Schmuland (University of Alberta) Rademacher’s theorem on configuration space

Qi-Man Shao (University of Oregon) Gaussian correlation conjecture and small ball probabilities

Zhan Shi (Université Paris VI) The maximum of the uniform emprical process

Gordon Slade (McMaster University) The scaling limit of the incipient infinite cluster in highdimensional percolation

Chistopher G. Small (University of Waterloo) The analysis of random shapes

Barbara Szyszkowicz (Carleton University) An Interplay of weighted approximations and change-point analysis

Keith Worsley (McGill University) The geometry of correlation fields, with an application to functional connectivity of the brain Hao Yu (University of Western Ontario) Weighted Kolmogorov-Smirnov test of stock return distributions

Ricardas Zitikis (Carleton University) The Vervaat, Lorenz and some other related processes of probability and mathematical statistics in weighted metrics

Topology - 4 Sub-sessions/ Topologie - 4

sous-sessions (Org: Eddy Campbell)

1) Differential Geometry and Global Analysis

Géométrie différentielle et analyse globale (Org: Muang Min-Oo, McKenzie Wang)

Igor Belagradek (McMaster University) Pinching and Pontrjagin classes

Christoph Böhm (McMaster University) Inhomogenous Einstein metrics on spheres

Jeffrey Boland (McMaster University) Magnetic fields on negatively curved manifolds

Virginie Charette (University of Maryland) Properly discontinous actions of free groups on Minkowski space

Jingyi Chen (University of British Columbia) Triholomorphic curves and complex ASD connections

Ailana Fraser (Stanford University) On the free boundary variational problem for minimal disks

Mike Gage (University of Rochester) Remarks on B. Süssman’s proof of the Banchoff-Pohl inequality

Miroslav Lovric (McMaster University) Multivariate normal distributions parametrized as a Riemannian symmetric space

Mohan Ramachandran (SUNY)

Patrick Ryan (McMaster University) Real hypersurfaces in complex space forms

Patrice Sawyer (Laurentian University) Ghostly symmetric spaces

Alina Stancu (Courant Institute) Asymptotoic behaviour of a crystalline evolution

John Toth (McGill University)

2) Homotopy Theory/Théorie de l’homotopie (Org: Lisa Langsetmo, Jim Shank)

Dan Christensen (Johns Hopkins University) Phantom maps: all or nothing

Fred Cohen (University of Rochester) On stunted projective spaces

Gustavo Granja (MIT) On self maps of HP^{n}

Steve Halperin (University of Toronto) The homotopy Lie algebra of a finite complex

Philip Heath (Memorial University) Fibre techniques in Nielsen perioidic point theory

Barry Jessup (University of Ottawa) Estimating the rational of LS-category of elliptic spaces

Brenda Johnson (Union College) Constructing and characterizing degree n functors

Keith Johnson (Dalhousie University) Elliptic homology cooperations

Sadok Kallel (University of British Columbia) The homology structure of free loop spaces

Kathryn Lesh (University of Toledo) Progress toward a partial splitting of E_{2} in the UASS for SO

Gaunce Lewis (Syracuse University) Recent results on Mackey functors for a compact Lie group

John Martino (University of Western Michigan) A Minami-Webb formula for compact Lie groups

Joe Neisendorfer (Rochester University) James-Hopf invariants, Anick’s spaces, and decompositions of the double loops on a Moore space

Stewart Priddy (Northwestern University) Decompsing products of classifying spaces

Charles Rezk (Northwestern University) A model for the homotopy theory of homotopy theory

Laura Scull (University of Chicago) Rational S^{1}- equivalent homotopy theory

Paul Selick (University of Toronto) Natural decompositions of loop suspensions and tensor algebras

Stephen Theriault (MIT) Homotopy exponents for certain mod-2(r) Moore Spaces 3) Set Theoretic Topology/Topologie des

ensembles (Org: Juris Steprans, Steve Watson)

Murray Bell (University of Manitoba) Cardinal functions of centered spaces

Maxim Burke (University of Prince Edward Island) Continuous functions which take a somewhere dense set of values on every open set

Krzysztof Cielsielski (West Virginia University) Each Polish space is cocompactly quasimetrizable

Wistar Comfort (Wesleyan University) Continuos cross sections on Abelian groups equipped with a Bohr Topology

Ilijas Farah (York University) Exp(N^{∗}) need not be an N^{∗}-image

Douglas Grant (University College of Cape Breton) Alternative universes: the role of set theory in topological algebra

Gary Gruenhage (Auburn University) More on alpha-Toronto spaces

Melvin Henriksen (Harvey Mudd College) Embedding a ring of continuous functions in a regular ring; preliminary report

James Hirschorn (University of Toronto) Towers of measurable functions

Valery Miskin (Kemerovo State University) Set ideals everywhere

Justin Moore (University of Toronto) A linearly fibered Souslinean space under Martin’s axiom

Eduardo Santillan (UNAM - Mexico) Topological properties of removable singularities for analytic functions

Alexander Shibakov (Tennessee Technological University) Controlling sequential order in topological vector spaces

Slawomir Solecki (Indiana University) Polish group actions and measures

Paul Szeptycki (Ohio University) Normality and property (a)

Andrzej Szymanski (Slippery Rock University of Pennylvania) On a class of special Namioka spaces

Franklin Tall (University of Toronto) The topology of elementary submodels Murat Tuncali (Nipissing University) On Generalizations of the Hahn-Mazurkiewicz Theorem

Ed Tymchatyn (University of Saskatchewan) Measures and topological dynamics on Menger manifolds

Grant Woods (University of Manitoba) Recent Developments in Rings of Continuous Functions

Eduardo Santillan (York University, Cinvestav, Mexico) Topological properties of removable singularities for analytic functions

4) Symplectic/Low Dimensional Topology Topologie en basses dimension/ Topologie symplectique (Org: Steve Boyer, Jacques Hurtubise, François Lalonde)

Steve Boyer (Université de Québec à Montréal) Norm duality and hyperbolic 3-manifolds

Jacques Hurtubise (McGill University, CRM) Framed parabolic bundles

François Lalonde (Université de Québec à Montréal, CRM) Applications of Quantum cohomology to some fundamental problems of dynamics

Yuriy Shkolnikov (University of Calgary) A generalisation of Whitney’s trick in dimension 4, borromeanism and related questions

Jennifer Slimowitz (Université de Québec à Montréal) Length minimizing geodesics in the group of hamiltonian diffeomorphisms

Xingriu Zhang (SUNY Buffalo) On simple points of the character variety of a cusped hyperbolic 3manifold

Universal Algebra and Multiple-Valued Logic Algèbre universelle et logique multiple valeurs (Org: L. Haddad)

Clifford Bergman (Iowa State University) Complexity of some Problems in Universal Algebra

Jie Fang (Simon Fraser University) Ockham algebras with pseudocomplementation

Ibrahim Garro (University of Toronto) An application of non-well founded sets to infinite valued infinitary propositional calculus George Grätzer (University of Manitoba) Independence Theorems for automorphism groups and congruence lattices of lattices

Jennifer Hyndman (University of Northern British Columbia) Dualizable is not the same as fully dualizable

Benoit Larose (Concordia University) Projective graphs and Hedetnyemi’s conjecture

Jonathan Leech (Westmount College) Noncommutative lattices: foundational issues and recent results

Robert Quackenbush (University of Manitoba) Varieties of binary linear codes

Ivo Rosenberg (Université de Montréal) Completeness for uniformly delayed circuits

Ross Willard (University of Waterloo) Independence of the linear commutator

CONTRIBUTED PAPERS COMMUNICATIONS LIBRES (Org: L. Haddad)

Alexandru Gica (University of Bucharest) A conjecture which implies the theorem of Gauss-Heegner-Stark-Baker

Igor Nikolaev (Université de Montréal) 3manifolds, foliations and K_{0}(C(X) ⋊ Z)

Dieter Ruoff (University of Regina) Solution of a Non-Euclidean convexity problem

Andrew Toms (Fields Institute, University ofToronto) Perforated K_{0}−groups of C^{∗}−algebras